# Spectral subspace extraction via incoherent quantum phase estimation

### v1

@@ IV Applications
In the following, we show three applications of DOS-QPE: the Fermi-Hubbard model, an electronic structure Hamiltonian, and a nuclear Hamiltonian. We use $\text{t}|\text{ket}\rangle$ [58] to construct the circuits, Qulacs for state vector evaluation [59], and OpenFermion for constructing qubit Hamiltonians [60]. To evaluate the molecular integrals in the electronic structure section, we use the PySCF Python package [61].
@@ IV.1 The Fermi-Hubbard Model
The Fermi-Hubbard model [62, 63, 64, 65] is one of the standard models of condensed matter physics as it describes relevant features of strongly-correlated electronic systems. It has also become a standard benchmark for Hamiltonian simulation algorithms [66, 67, 68] in the quantum computing community since exact solutions can be found for the one-dimensional case [69]. The study of the two- and higher-dimensional cases is a topic of intense research activity, where only numerical simulations are available [70, 71, 72]. Its Hamiltonian is written as
@@
In the remainder of this section, we apply the DOS-QPE algorithm to the Fermi-Hubbard chain model with three sites. We choose $t=1$ and $U=4$ such that we are in the so-called “intermediate coupling” regime, under which it is difficult classically to extract even ground-state properties from numerical simulations at half-filling [73]. We can then map this second-quantized Hamiltonian, using the Jordan-Wigner encoding, to a linear combination of products $P$ of Pauli operators $\sigma_{j}$,
@@ IV.2 Electronic Structure
respectively, and are routinely obtained as part of a classically inexpensive Hartree-Fock calculation. The quantities $\phi_{p}$ are the molecular orbitals and form the 1-particle wavefunctions mentioned in Sec. II.2, $\mathbf{r}$ are the positions of the electrons, and $Z_{A}$ and $\mathbf{R_{A}}$ are the charges and positions of the nuclei, respectively. For the same reasons as in Sec. II.2 and the former section, the Jordan-Wigner mapping is selected.
@@
We choose the twisted conformation of ethylene as the example chemical system, since by varying the pyramidalization angle we expect to encounter a conical intersection between the two lowest lying singlet eigenstates, and a crossing of the lowest lying singlet state with the triplet ground state [74]. We first optimise the planar geometry at the CCSD(T) level classically in the cc-pVDZ basis, then rotate the two hydrogens on one carbon atom to have dihedral angles with the hydrogen atoms on the other carbon atom of exactly $90^{\circ}$. We then vary the angle $\alpha$ as shown in Fig. 3a). We select a 2-electrons in 2-orbitals active space, and along the reaction coordinate, evaluate the overlap of the orbitals at each point with the orbitals from the previous point, such that we can place the orbitals at maximum coincidence at each step along the potential energy surfaces and ensure a consistent orbital selection. Such that the eigenvalues lay in a fixed interval and the movement of each peak in Fig. 3c–f) is smooth, and to obtain a smooth curve in Fig. 3b), we find our target Hamiltonians with the following formula,
@@ IV.3 Nuclear Structure
The description of nuclear structure in atomic nuclei follows a similar prescription as that of molecules in quantum chemistry within the nuclear shell model [75]. Under this approach, protons and neutrons (collectively called nucleons), basic fermion components of a nucleus, move in a set of single-particle states that form a valence space. These single-particle states, also called orbitals, are labeled by the quantum numbers $(n,l,j,m,t_{z})$, where $n$ is the principal quantum number, $l$ is the orbital angular momentum, $j$ is the total angular momentum after coupling with the nucleon spin12-$\frac${1}{2}, $m$ is its third-component projection and $t_{z}$ is the third-component projection of the isospin $t=\frac{1}{2}$, an additional quantity equivalent to the spin defined in order to discern between protons and neutrons.
The nuclear shell model can be described by means of a second-quantized Hamiltonian as in Eq. (23), where the single-particle energies $h_{pp}$ and two-body matrix elements $h_{pqrs}$ are adjusted to reproduce observable properties of selected nuclei [76]. The nuclear states are described by a well-defined total angular momentum $J$ and isospin $T$, resulting from the coupling of individual nucleonic $j$ and $t$; with their corresponding third-component projections $M$ and $T_{z}$, which are equal to the sum of the third components of the nucleons in the valence space, $m$ and $t_{z}$. Thus, a many-body basis state can be written in terms of states with well-defined $M$ and $T_{z}$
@@
Many quantum algorithms have been applied to obtain nuclear-structure properties based on the ground state [77, 78, 79, 80, 81, 82, 83, 84] and also low-lying excited states [85, 86, 87, 88]. The structure and entanglement features of nuclear states have also been thoroughly studied [89, 90, 91, 92].
Figure 4: a) Many-body state basis used for 6Be. Label represents the indices in the corresponding Hamiltonian matrix in this many-body basis. The many-body states are composed of single-particle states from all possible combinations of two valence protons in the $p$-shell variational space of panel b) satisfying the condition $M=m_{1}+m_{2}=0$. b) Configurational valence space for the $p$-shell. The valence space is decomposed in two degenerate $j$ orbitals with degeneracy $(2j+1)$. The labels on the left correspond to the standard spectroscopic notation $n\ell_{j}$, where $\ell=p$ means $\ell=1$. Numbers on top of every single-particle state represent the labels used in this work. Slightly modified from [80].
In this work, we aim to resolve the full spectrum in a single calculation with DOS-QPE. We simulate the 6Be nucleus, which is composed of two protons in the valence space of a p shell, as shown in Fig. 4, with six single-particle states for the protons. Thus, we have in this case two identical particles moving in the valence space, and consequently, the total isospin is $T=1$. The Hamiltonian used is based on the phenomenological Cohen-Kurath interaction [93].
While we could directly transform the second-quantized Hamiltonian using the Jordan-Wigner mapping as demonstrated in the former examples, for this specific case we opt to build the Hamiltonian matrix using the minimal many-body basis spanned for 6Be, consisting of only five states whose total angular momentum projection $M$ is equal to zero [80]. The many-body basis used for 6Be is listed in the table of Fig. 4a). The resulting Hamiltonian matrix is then mapped onto the qubit operator basis using the Linear Combination of Unitaries (LCU) formalism [94]. Thus, using this mapping, we managed to reduce the size of the Hamiltonian from six to three qubits, at the expense of three additional degenerate eigenvalues equal to zero.
@@
In light of these results, although demonstrated on a small example, we may conclude that DOS-QPE could become a promising tool for efficiently extracting spectral information and thus dynamical processes [95, 96, 97, 98, 99] in atomic nuclei for near-term quantum devices.
@@ V Discussion and conclusion
We have shown how, by using a generalization of quantum phase estimation probed with maximally mixed or symmetry-adapted ensemble states, we can access spectral information of the generating Hamiltonians. By embedding QPE within a purified framework [4], the traditional eigenstate-dependent algorithm transforms into a flexible tool for ensemble-based spectral analysis. This approach grants access to the density of states of Hamiltonians, effectively revealing their spectral properties. From here, thermodynamic and dynamical properties of systems become accessible while mitigating the effects of noisy or imperfect state preparations.
@@
We framed the recovery of discrete spectral features from the output of DOS-QPE as a convex optimization problem. We resolve spectral structures exploiting sparsity and physical constraints to produce high-fidelity reconstructions. We propose a modular postprocessing that restores physically meaningful integer-value spectra.
@@
While we have demonstrated DOS-QPE as a flexible primitive for ensemble-based quantum spectroscopy across fermionic and nuclear models, optimizations or resource reductions might be desirable. One immediate goal is to reduce circuit complexity and resource cost, tailoring the primitive for currently available quantum hardware [100]. Implementing single-ancilla variants [101, 39, 102] or taking advantage of shadow techniques [103, 104, 105] could significantly lower circuit depth and overhead, making the primitive viable for near-term applications.
@@
Further directions include exploring alternative probe states, such as spin-selective or symmetry constrained ensembles, which may target subspace-resolved spectral features often inaccessible to conventional methods. In this context, developing spectrum manipulation techniques to enhance signal contrast, reduce costs, and mitigate noise could be particularly valuable. Such strategies would accelerate the reconstruction process and alleviate the current bottlenecks in measurement and post-processing.
@@
DOS-QPE shows potential as a framework that bypasses state preparation bottlenecks while taking advantage of the incoherent nature of the environment to return physically relevant observables otherwise inaccessible. Connections to non-linear functions of spectral densities, gradient estimation, and Bell-type measurements [106] indicate a broader algorithmic utility within hybrid optimization and quantum learning frameworks. Embedding DOS-QPE within Green’s functions formulations [107, 108] could further help in dynamical response theory offering access to electronic, vibrational, or nuclear observables.

### v2

@@ IV Applications
In the following, we show three applications of DOS-QPE: the Fermi-Hubbard model, an electronic structure Hamiltonian, and a nuclear Hamiltonian. We use $\text{t}|\text{ket}\rangle$ [sivarajah2020t] to construct the circuits, Qulacs for state vector evaluation [suzuki2021qulacs], and OpenFermion for constructing qubit Hamiltonians [mcclean2020openfermion]. To evaluate the molecular integrals in the electronic structure section, we use the PySCF Python package [sun2018pyscf].
@@ IV.1 The Fermi-Hubbard Model
The Fermi-Hubbard model [hubbardI, hubbardII, hubbardIII, hubbardIIII] is one of the standard models of condensed matter physics as it describes relevant features of strongly-correlated electronic systems. It has also become a standard benchmark for Hamiltonian simulation algorithms [kan2024resource, babbush2018encoding, arute2020observation] in the quantum computing community since exact solutions can be found for the one-dimensional case [lieb1968absence]. The study of the two- and higher-dimensional cases is a topic of intense research activity, where only numerical simulations are available [chen2021quantum, varney2009quantum, pasqualetti2024equation]. Its Hamiltonian is written as
@@
In the remainder of this section, we apply the DOS-QPE algorithm to the Fermi-Hubbard chain model with three sites. We choose $t=1$ and $U=4$ such that we are in the so-called “intermediate coupling” regime, under which it is difficult classically to extract even ground-state properties from numerical simulations at half-filling [fradkin2015colloquium]. We can then map this second-quantized Hamiltonian, using the Jordan-Wigner encoding, to a linear combination of products $P$ of Pauli operators $\sigma_{j}$,
@@ IV.2 Electronic Structure
respectively, and are routinely obtained as part of a classically inexpensive Hartree-Fock calculation. The quantities $\phi_{p}$ are the molecular orbitals and form the 1-particle wavefunctions mentioned in Sec. II.1.2, $\mathbf{r}$ are the positions of the electrons, and $Z_{A}$ and $\mathbf{R_{A}}$ are the charges and positions of the nuclei, respectively. For the same reasons as in Sec. II.1.2 and the former section, the Jordan-Wigner mapping is selected.
@@
We choose the twisted conformation of ethylene as the example chemical system, since by varying the pyramidalization angle we expect to encounter a conical intersection between the two lowest lying singlet eigenstates, and a crossing of the lowest lying singlet state with the triplet ground state [xu2025conical]. We first optimise the planar geometry at the CCSD(T) level classically in the cc-pVDZ basis, then rotate the two hydrogens on one carbon atom to have dihedral angles with the hydrogen atoms on the other carbon atom of exactly $90^{\circ}$. We then vary the angle $\alpha$ as shown in Fig. 3a). We select a 2-electrons in 2-orbitals active space, and along the reaction coordinate, evaluate the overlap of the orbitals at each point with the orbitals from the previous point, such that we can place the orbitals at maximum coincidence at each step along the potential energy surfaces and ensure a consistent orbital selection. Such that the eigenvalues lay in a fixed interval and the movement of each peak in Fig. 3c–f) is smooth, and to obtain a smooth curve in Fig. 3b), we find our target Hamiltonians with the following formula,
@@ IV.3 Nuclear Structure
The description of nuclear structure in atomic nuclei follows a similar prescription as that of molecules in quantum chemistry within the nuclear shell model [caurier2005shell]. Under this approach, protons and neutrons (collectively called nucleons), basic fermion components of a nucleus, move in a set of single-particle states that form a valence space. These single-particle states, also called orbitals, are labeled by the quantum numbers $(n,l,j,m,t_{z})$, where $n$ is the principal quantum number, $l$ is the orbital angular momentum, $j$ is the total angular momentum after coupling with the nucleon spin12-$\frac${1}{2}, $m$ is its third-component projection and $t_{z}$ is the third-component projection of the isospin $t=\frac{1}{2}$, an additional quantity equivalent to the spin defined in order to discern between protons and neutrons.
The nuclear shell model can be described by means of a second-quantized Hamiltonian as in Eq. (27), where the single-particle energies $h_{pp}$ and two-body matrix elements $h_{pqrs}$ are adjusted to reproduce observable properties of selected nuclei [poves1981theoretical]. The nuclear states are described by a well-defined total angular momentum $J$ and isospin $T$, resulting from the coupling of individual nucleonic $j$ and $t$; with their corresponding third-component projections $M$ and $T_{z}$, which are equal to the sum of the third components of the nucleons in the valence space, $m$ and $t_{z}$. Thus, a many-body basis state can be written in terms of states with well-defined $M$ and $T_{z}$
@@
Many quantum algorithms have been applied to obtain nuclear-structure properties based on the ground state [kiss2022quantum, dumitrescu2018cloud, romero2022solving, perez2023nuclear, stetcu2022variational, costa2024quantum, sarma2023prediction, carrasco2025comparison] and also low-lying excited states [zhang2025excited, bhoy2024shell, ruiz2022accessing, robin2023quantum]. The structure and entanglement features of nuclear states have also been thoroughly studied [robin2021entanglement, hengstenberg2023multi, perez2023quantum, tichai2023combining].
Figure 4: a) Many-body state basis used for 6Be. Label represents the indices in the corresponding Hamiltonian matrix in this many-body basis. The many-body states are composed of single-particle states from all possible combinations of two valence protons in the $p$-shell variational space of panel b) satisfying the condition $M=m_{1}+m_{2}=0$. b) Configurational valence space for the $p$-shell. The valence space is decomposed in two degenerate $j$ orbitals with degeneracy $(2j+1)$. The labels on the left correspond to the standard spectroscopic notation $n\ell_{j}$, where $\ell=p$ means $\ell=1$. Numbers on top of every single-particle state represent the labels used in this work. Slightly modified from [perez2023nuclear].
In this work, we aim to resolve the full spectrum in a single calculation with DOS-QPE. We simulate the 6Be nucleus, which is composed of two protons in the valence space of a p shell, as shown in Fig. 4, with six single-particle states for the protons. Thus, we have in this case two identical particles moving in the valence space, and consequently, the total isospin is $T=1$. The Hamiltonian used is based on the phenomenological Cohen-Kurath interaction [Cohen_Kurath].
While we could directly transform the second-quantized Hamiltonian using the Jordan-Wigner mapping as demonstrated in the former examples, for this specific case we opt to build the Hamiltonian matrix using the minimal many-body basis spanned for 6Be, consisting of only five states whose total angular momentum projection $M$ is equal to zero [perez2023nuclear]. The many-body basis used for 6Be is listed in the table of Fig. 4a). The resulting Hamiltonian matrix is then mapped onto the qubit operator basis using the Linear Combination of Unitaries (LCU) formalism [childs2012hamiltonian]. Thus, using this mapping, we managed to reduce the size of the Hamiltonian from six to three qubits, at the expense of three additional degenerate eigenvalues equal to zero.
@@
In light of these results, although demonstrated on a small example, we may conclude that DOS-QPE could become a promising tool for efficiently extracting spectral information and thus dynamical processes [roggero2020preparation, stetcu2025quantum, garcia2024nuclear, zhang2025quantum, du2025systematic] in atomic nuclei for near-term quantum devices.
@@ V Discussion and conclusion
We have shown how, by using a generalization of quantum phase estimation probed with maximally mixed or symmetry-adapted ensemble states, we can access spectral information of the generating Hamiltonians. By embedding QPE within a purified framework [scali2025signed], the traditional eigenstate-dependent algorithm transforms into a flexible tool for ensemble-based spectral analysis. This approach grants access to the density of states of Hamiltonians, effectively revealing their spectral properties. From here, thermodynamic and dynamical properties of systems become accessible while mitigating the effects of noisy or imperfect state preparations.
@@
We framed the recovery of discrete spectral features from the output of DOS-QPE as a convex optimization problem. Exploiting sparsity and physical constraints, we resolve spectral structures and obtain high-fidelity reconstructions, and we propose a modular postprocessing pipeline that restores physically meaningful integer-valued spectra. Together with the circuit primitive, this yields an end-to-end workflow from mixed-state probes to reconstructed eigenspectra.
@@
While we have demonstrated DOS-QPE as a flexible primitive for ensemble-based quantum spectroscopy across fermionic and nuclear models, further optimizations and resource reductions are desirable. One immediate goal is to reduce circuit complexity and resource cost, tailoring the primitive for currently available quantum hardware [fujitsu_qpu]. Implementing single-ancilla variants [Ding_2023, Ding_2024_2, Ding_2024_3] or taking advantage of shadow techniques [Huang_2020, Elben_2022, Faehrmann_2025] could significantly lower circuit depth and overhead, making the primitive viable for near-term applications.
@@
Additional directions include exploring alternative probe states, such as spin-selective or other symmetry-constrained ensembles, which may target subspace-resolved spectral features often inaccessible to conventional methods. In this context, developing spectrum manipulation techniques to enhance signal contrast, reduce costs, and mitigate noise could be particularly valuable, accelerating the reconstruction process and alleviating current bottlenecks in measurement and post-processing.
@@
DOS-QPE thus emerges as a framework that bypasses eigenstate-preparation bottlenecks while leveraging the incoherent nature of realistic environments to return physically relevant observables that are otherwise difficult to access. Conceptually, this purified viewpoint resonates with the random purification channel paradigm developed in quantum learning theory by Tang et al. [tang], Pelecanos et al. [pelecanos], and Girardi et al. [girardi]: in both cases, purification elevates mixed-state access from a nuisance to a resource for running pure-state algorithms. Connections to non-linear functions of spectral densities, gradient estimation, and Bell-type measurements [Scali_2024_2] indicate a broader algorithmic utility within hybrid optimization and quantum learning frameworks. Embedding DOS-QPE within Green’s function formulations [Libbi_2022, Umeano_2025] could further support dynamical response theory, offering access to electronic, vibrational, or nuclear observables and strengthening the role of purified phase estimation as a core primitive for future quantum simulations.
Author contributions
S.S. and J.K. had the idea and developed the methods of the paper. J.K. wrote Dicke State, FH and Electronic Structure sections. A.M.R. developed the nuclear application and wrote the relative section. S.S. linked the paper to the Random Purification Channel, wrote the rest of the manuscript including appendices and prepared all figures. All authors reviewed the manuscript.
Funding
This work has been entirely funded by Fujitsu Research of Europe.